Operations Report – February 11th

Crew 240 Operations Report 11-02-2022

SOL: 11

Name of person filing report: François Vinet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: yes

Hours: 154.8

Beginning charge: 100%

Ending charge: 71%

Currently charging: left uncharged, handled by Atila

Opportunity rover used: no

Hours: 83.7

Beginning charge: NA

Ending charge: NA

Currently charging: handled by Atila

Curiosity rover used: yes

Hours: 166.5

Beginning charge: 100%

Ending charge: 67%

Currently charging: left uncharged, handled by Atila

Perseverance rover used: no

Hours: 227.0

Beginning charge: NA

Ending charge: NA

Currently charging: handled by Atila

General notes and comments: NA

Summary of Hab operations:

WATER USE: 28 gallons

Water (static tank): 120 gallons

Water (loft tank): 42 gallons

Water Meter: 0155722.0 units

Static to Loft Pump used – yes

Static tank pipe heater (on or off): on

Static tank heater (On or off): on

Toilet tank emptied: yes

Summary of internet: NA

Summary of suits and radios: NA

Summary of GreenHab operations:

WATER USE: 10 gallons at 8:10am and 8 gallons at 6pm

Heater: On

Supplemental light: On

Harvest: 90g Cherry tomatoes

Note: Water level in the GreenHab water tank is getting low, instead of using the top tap, I am now using the bottom one to fill up the watering can.

Summary of ScienceDome operations:

Dual split: off

SOC: 75% at 7.10pm

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: NA

Questions, concerns and requests to Mission Support: NA

EVA Report – February 11th

Crew 240 EVA Report 08-02-2022

EVA # 12

Author: Julie Levita

Purpose of EVA: Exploration and looking out for interesting features for 3D mapping. Changing batteries on atmospheric experiment

Start time: 9:40

End time: 12:05

Narrative: Everything went great! The landscape was beautiful and very interesting. We think it is a good spot to try some 3D mapping with the drone. We were able to change out the battery for the atmospheric experiment without any problem.

Destination: Lillith Canyon, Exploration zone D (through Special Region, Exploration Zone C)

Coordinates (use UTM NAD27 CONUS): 518 000, 425 600

Participants: Julie Levita, Clément Plagne, Francois Vinet, Maxime Jalabert

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Rovers, foot

Sol Summary – February 11th

Crew 240 Sol Summary Report 11Feb2022

Sol: 11

Summary Title: Tirelessness

Author’s name: Clément Plagne

Mission Status: Nominal

Sol Activity Summary: Had this been a standard, two-week rotation like the one I did the first time I was at MDRS, today would basically be the time for a farewell to sim, and for planning for getting back onto good old Earth. Seen from this angle, it’s clear that I want to try to get as much done in these few remaining days as I can, while wrestling with the fatigue of the past two weeks.

There’s still a lot to work on on the atmospheric experiment side – some instruments had issues from the trip, and fixing them took a while, but it’s looking like we’ll get at least a solid week of data before handing them over to the next crew, also from Supaero. And there’s still some places to get the drones running on, and 3D maps to be made.

Meanwhile, we went to the Special Region today – far further North than I did two years ago, and there we saw a diversity of landscapes and terrain that will probably prove fascinating on Pierre’s drone shots, and, if he manages to process them as well, 3D maps. The afternoon was spent finishing up on the human research experiments, and while we’re looking forward to a slightly busy weekend, it’ll still be a time of calm in comparison.

Look Ahead Plan: Finishing up work on the atmospheric experiments we still haven’t installed yet, catching a little rest.

Anomalies in work: None to report

Weather: Shockingly warm

Crew Physical Status: All are well

EVA: One, carried out nominally

Reports to be filed: Astronomy Report, Sol Summary, HSO Report, Operations Report, Journalist Report, Crew Pictures, EVA Request, EVA Report

Support Requested: None in particular.

Copyright © The Mars Society. All rights reserved. | Main Site